I need to solve using substitution:
2x+5y =25
3x-2y+4z=1
4x-3y+ z=9

Please help I am so confused!
Thanks,
Cecilia

Answer by venugopalramana(3286)

2x+5y =25....................A
3x-2y+4z=1......................B
4x-3y+ z=9 ............................C
LET US ELIMINATE Z USING B & C EQNS.WE FIND Z FROM EQN.C AND SUBSTITUTE IT IN EQN.B.
FROM C .....
Z=9-4X+3Y...................D
SUBSTITUTING IN B WE GET
3X-2Y+4(9-4X+3Y)=1
3X-2Y+36-16X+12Y=1
-13X+10Y=-35....................E
MULTIPLY EQN A WITH 2 AND SUBTRACT FROM EQN.E
2*A.............4X+10Y=50..........................F
EQN E - EQN. F .................
-13X+10Y-4X-10Y=-35-50=-85
-17X=-85
17X=85
X=85/17=5
NOW PUTTING IN A
2*5+5Y=25
5Y=25-10=15
Y=15/5=3
NOW PUTTING IN C..................
4*5-3*3+Z=9
Z=9+9-20=-2
SO X=5,Y=3 AND Z=-2